OpenStudy (anonymous):

how do I know that 5/12 is less than 1/2 ?

OpenStudy (anonymous):

6/12 = 1/2, therefore 5/12 < 1/2

OpenStudy (raffle_snaffle):

long divison

OpenStudy (raffle_snaffle):

or use calculator

OpenStudy (anonymous):

but how do I know

OpenStudy (anonymous):

Do you know the least common denominator?

OpenStudy (anonymous):

no

OpenStudy (usukidoll):

decimal values?

OpenStudy (anonymous):

Okay so start with 1/2 The denominator (bottom) is 2 Multiply top and bottom of 5/12 by 2 and get 10/24 Now start with 5/12 The denominator is 12 Multiply top and bottom of 1/2 by 12 and get 12/24 10/24 < 12/24

OpenStudy (anonymous):

Once the denominator is the same, then the numerator will determine which is bigger.

OpenStudy (anonymous):

1/2 = 0.500, 5/12 = 0.416667

OpenStudy (anonymous):

ohhhhhhhhhhhhh ok thanks SO MUCH!!!!!!!!!!!